We unpacked our MIT instrument and set it all up. It worked fine except for the computer networking (so our computer talks to the other computers here and at MIT). We are up here now (in the morning) to try to get the networking going. It's been several hours now and the network is not working. Jeff is on the phone with Eric, a computer expert over in Honolulu, trying to get it fixed. Hey good news: they just fixed it. The bad news is that the IRTF camera had been turned off before the last crew abandoned the summit due to the blizzard, and the detector was warm (normally it is kept extremely cold to make it work better). One of the telescope staffers is working on the IRTF camera, to see if we can get it cold enough to work tonight. It seems grim. This is not a "show-stopper" however, because our MIT camera is working.
The REALLY bad news, though, is that the entire dome is encased in thick ice - at some points over a foot thick! This is very bad because you cannot open the dome - big chunks of ice might fall in and damage the telescope. You also cannot rotate the dome, since the ice is sealing it shut. The crew is outside assessing the situation, but they think it is unlikely we will be able to get the dome open tonight. I can hear them banging on the dome right now, even as I am typing. They are trying to break loose some of the ice. Last week, some crew members from another telescope were injured trying to remove ice from a dome, so everyone here is very cautious. The winds are blowing over 60 mph, and the temperature here is hovering at 30 degrees - so the ice may not melt in the sunshine. This ice could be the show-stopper.
The skies are absolutely crystal clear - a gorgeous deep blue, with not a cloud in sight. It is now 17 hours until the occultation.